# VIN-NBV: A View Introspection Network for Next-Best-View Selection for Resource-Efficient 3D Reconstruction

### v1

@@ Abstract
Next Best View (NBV) algorithms aim to acquire an optimal set of images using minimal resources, time, or number of captures to enable efficient 3D reconstruction of a scene. Existing approaches often rely on prior scene knowledge or additional image captures and often develop policies that maximize coverage. Yet, for many real scenes with complex geometry and self-occlusions, coverage maximization does not lead to better reconstruction quality directly. In this paper, we propose the View Introspection Network (VIN), which is trained to predict the reconstruction quality improvement of views directly, and the VIN-NBV policy. A greedy sequential sampling-based policy, where at each acquisition step, we sample multiple query views and choose the one with the highest VIN predicted improvement score. We design the VIN to perform 3D-aware featurization of the reconstruction built from prior acquisitions, and for each query view create a feature that can be decoded into an improvement score. We then train the VIN using imitation learning to predict the reconstruction improvement score. We show that VIN-NBV improves reconstruction quality by ${\sim}30\%$ over a coverage maximization baseline when operating with constraints on the number of acquisitions or the time in motion.
Figure 1: We present VIN-NBV, an NBV policy that selects Next Best Views by maximizing predicted reconstruction quality with limited resources. VIN-NBV (blue) outperforms coverage-based Cov-NBV (red), achieving higher reconstruction accuracy with fewer acquisitions.
Keywords: 3D Reconstruction, Next Best View, Imitation Learning
22footnotetext: Project page: Here
@@ 1 Introduction
Acquiring 3D knowledge of an environment is often a crucial step for many robotics applications; e.g., a drone surveying a disaster zone to gather information for assisting search and rescue efforts, or autonomous robots monitoring large construction and agricultural sites to gather necessary information. Real-world environments contain diverse objects of varying sizes, occluding objects, and areas with important fine-grained details; such complexities currently require a slow dense scan [1, 2, 3, 4, 5, 6, 7, 8] to effectively reconstruct the 3D scene. However, in many of these scenarios, acquiring the 3D scene in the shortest time possible is often critical, e.g. for search and rescue, time is of essence, while for monitoring large construction or agricultural sites, battery life is often limiting. This often led researchers to develop techniques that are able to more judiciously scan a novel environment for resource-efficient 3D reconstruction, where resource constraints can be expressed as the number of acquisitions, distance traversed, acquisition time, or battery life. This problem has often been called the Next Best View (NBV) selection problem, where the goal is to predict a set of optimal camera viewpoints that can maximize reconstruction quality.
@@
Predicting NBVs is a challenging problem due to the large search space for selecting the pose of each view, the challenge in efficiently computing the optimal solution in this non-convex space, and the complexity of scene geometry and occlusions. Earlier works on NBV often assume prior knowledge of the scene (e.g., a preliminary scan or CAD model) [9, 10, 11, 12, 13, 14], which limits applications in previously unexplored environments. Approaches that do not require prior knowledge of the scene predict NBVs by either maximizing coverage in the scene [15, 16, 17, 18, 19, 20] or by maximizing information gain [21, 11, 22] in the image space. While earlier methods have relied on heuristics or optimization [15, 16], recent methods often train a deep reinforcement learning (RL) algorithm [17, 18, 21, 11, 23, 22, 24, 25] to maximize coverage.
@@
While maximizing coverage with RL [25, 18] may lead to a generalizable policy that can provide a good approximation of the scene, it ignores the fact that certain regions in a scene have more complicated geometry and self-occlusion than others. This means certain regions in the scene might require more captures than other regions to generate higher-quality reconstruction. For example, if the fence of the house occludes a part of the wall, a single viewpoint capture may satisfy the coverage criterion, but will lead to poor reconstruction with holes in the wall unless additional viewpoints are captured. Thus, our key idea is to develop an NBV acquisition policy that is trained to directly maximize the 3D reconstruction accuracy instead of only relying on the coverage criterion. We assume no prior information about the scene, and that the robot is equipped with an RGBD camera.
Our key idea is to formulate a greedy sequential next best view selection strategy, where at each step, we only capture a single ‘best’ next view that maximizes reconstruction improvement over the already captured set of base images. We repeat the process until some termination criterion has been reached. We achieve this by designing a View Introspection Network (VIN) that has been trained to assess the effectiveness of any query viewpoint in reducing reconstruction error over the observed base acquisitions. At each step of the acquisition, we uniformly sample query views around the object, evaluate their effectiveness by using the VIN, and choose the one that will maximize reconstruction improvement over the base views. We observe the greedy imitation learning approach of VIN-NBV to be more stable and effective than non-greedy RL based approaches.
Our key contribution is to design a lightweight neural network, VIN, that uses smart 3D-aware featurization to assess the effectiveness of any query viewpoint in reducing reconstruction error over an already captured set of views. We train VIN on simulated data using imitation learning, where we first calculate the true fitness score of the query view using relative reconstruction improvement over the base views. Our proposed VIN-NBV policy is generalizable and can operate under any kind of motion planning constraints and budget.
Experimental evaluation of the VIN-NBV policy with constraints on the number of acquisitions and time in motion indicate significant improvements of up to ${\sim}30\%$ over a coverage only baseline, especially during earlier stages of acquisition. Ultimately both methods converge to the same solution under a very large time or acquisition budget.
We also observe that VIN-NBV significantly outperforms existing RL-based algorithms that aim to maximize coverage. VIN-NBV reduces reconstruction error by 41% compared to Scan-RL [26] 39% compared to GenNBV [25].
@@ 2 Related work
NBV methods that require prior knowledge. Many existing approaches assume prior knowledge of the 3D scene, which limits applications in previously unexplored environments. Some approaches [9, 10, 11] directly exploit a pre-existing 3D model or its approximation. Some others rely on 2D maps, such as estimating the height of buildings as a rough 3D model [27] or obtaining a 2.5D model of the scene [14]. For scenes lacking pre-existing information, often a drone fly-through along a default trajectory is made to obtain an initial coarse reconstruction [16, 21]. In contrast, our method does not require any prior information about the scene and starts by acquiring two adjacent views and sequentially chooses the best next views until a specific termination criterion is met.
Optimal view selection from pre-capture imagery. Another line of work focuses on selecting an optimal subset of images from densely captured imagery to improve reconstruction quality. While earlier versions of these algorithms focused on selecting optimal views for Multiview Stereo Reconstruction [28, 1], more recent versions explored view selection for Neural Radiance Field (NeRF) and its variants [29, 30, 31, 22]. Our work is fundamentally different from this line of work, as we do not pre-capture the scene densely; rather, we only acquire images from the optimal viewpoints predicted by the NBV policy.
Criterion for predicting NBV. Previous approaches use proxy metrics like coverage [15, 18, 32, 17, 25] or information gain [33, 34, 35, 36] for predicting NBVs. State-of-the-art NBV techniques, GenNBV [25] and ScanRL [18], use Reinforcement Learning (RL) with coverage-based reward functions. While coverage-based policies may lead to reasonable reconstruction quality, they fail to account for complex structures and occlusions, often resulting in holes in the reconstructed scene. On the other hand, the information gain criterion is image-based and often lacks 3D knowledge of the scene, restricting generalization. Our work goes beyond coverage by directly maximizing the reconstruction quality, which leads to significant improvement under resource constraints, as shown empirically. Rather than relying on RL [18, 25], we design a greedy sequential policy, VIN-NBV, trained with imitation learning to help predict the reconstruction improvement of next views.
@@ 3 Method
We will first provide a mathematical overview of the NBV problem in Sec. 3.1. Then, we will introduce our proposed sampling-based greedy NBV policy, VIN-NBV, in Sec. 3.2, followed by the design of the View Introspection Network (VIN) in 3.3 and its training in 3.4.
@@ 3.1 Problem Setup
We consider a robot that has acquired $k$ initial base images of a scene $I_{base}=\{I_{1},...,I_{k}\}$ from viewpoints with camera parameters $C_{base}=\{C_{1},...,C_{k}\}$ and depth maps $D_{base}=\{D_{1},...,D_{k}\}$ either captured with a depth sensor or predicted with a Monocular Depth Estimator or Multi-View Stereo algorithms. The robot then runs a 3D reconstruction pipeline using the initial base views to reconstruct the 3D scene as $\mathcal{R}_{base}$. The goal of NBV is to predict a set of $m$ next best views $C_{nbv}=\{C_{{1}},...,C_{m}\}$ from which the scene should be captured to maximize the reconstruction quality of the scene.
More specifically, from $C_{nbv}$ we capture NBV images $I_{nbv}=\{I_{1},...,I_{m}\}$ and associated depth maps $D_{nbv}=\{D_{1},...,D_{m}\}$ and perform 3D reconstruction to create $\mathcal{R}_{final}$ using $I_{base}\cup I_{nbv}$ and $D_{base}\cup D_{nbv}$. While previous NBV techniques maximize coverage, we maximize reconstruction quality measured by the relative improvement of the chamfer distance of $\mathcal{R}_{final}$ over $\mathcal{R}_{base}$:
$C^{*}_{nbv}=\operatorname*{arg\,max}_{C_{nbv}}~{}\frac{|CD(\mathcal{R}_{final},\mathcal{R}_{GT})-CD(\mathcal{R}_{base},\mathcal{R}_{GT})|}{CD(\mathcal{R}_{base},\mathcal{R}_{GT})},$ (1)
@@
3.2 NBV Policy with View Introspection Network
{wrapfloat}
algorithmr0.53 VIN-NBV Policy
1:$I^{k}_{\mathrm{base}}\leftarrow\{I_{1},\dots,I_{k}\}$
2:$C^{k}_{\mathrm{base}}\leftarrow\{C_{1},\dots,C_{k}\}$
3:$D^{k}_{\mathrm{base}}\leftarrow\{D_{1},\dots,D_{k}\}$
@@ 3.2 NBV Policy with View Introspection Network
9:         ^$\widehat{\mathcal{RRI}}(q_{i})=\text{VIN}_{\theta}(R^{t}_{base},C^{t}_{base},C_{q_{i}})$
@@
We propose a simple greedy imitation learning based approach, where for each acquisition, the robot samples a set of query camera viewpoints and evaluates their fitness, and chooses the ‘best’ one and repeats the process until some termination criterion has been reached. This acquisition policy has been described in Algorithm 3.2.
@@
The VIN-NBV policy begins with $I^{k}_{base}$, $C^{k}_{base}$, $D^{k}_{base}$ then iteratively selects next views until we have reached a desired termination criteria defined to fit downstream task constraints, e.g., number of image acquisitions, time traversed, battery life, etc. During each step, we create the most recent version of our reconstruction $R^{t}_{base}$ by back-projecting our RGBD capture into 3D space. If RGB capture is made, 3D reconstruction techniques like Multiview Stereo or Monocular Depth Estimation can be used to reconstruct $R^{t}_{base}$. We then sample $n$ query views around this reconstruction to evaluate their fitness and choose the ‘best’.
@@
Our key idea is the introduction of the View Introspection Network (VIN), which independently evaluates potential query/next views and predicts their fitness. Instead of evaluating the fitness of each view to maximize coverage, we focus on maximizing the reconstruction quality. More specifically, we define fitness criterion as the Relative Reconstruction Improvement ($\mathcal{RRI}$) over the base views by capturing the query view $q$ as defined in eq. 2. Relative Improvement is formulated to be independent of object types and scales, which otherwise affects the Chamfer Distance computation.
$\mathcal{RRI}(q_{i})=\frac{|CD(\mathcal{R}_{base\cup q_{i}},\mathcal{R}_{GT})-CD(\mathcal{R}_{base},\mathcal{R}_{GT})|}{CD(\mathcal{R}_{base},\mathcal{R}_{GT})}.$ (2)
@@
We train VIN to predict RRI fitness criterion $\widehat{RRI}(q)$ for a query view $q$ by taking the existing reconstruction $R_{base}$, camera parameters $C_{base}$, and query view camera parameters $C_{q}$ as input:
^$\widehat{\mathcal{RRI}}(q)=\text{VIN}_{\theta}(R_{base},C_{base},C_{q}).$ (3)
@@
The design of VIN’s neural architecture and its training are described in Sec 3.3 and 3.4. After evaluating all $n$ query views, VIN-NBV greedily selects the one with the highest improvement score. We move to the selected view position $C^{t}_{*}$ and acquire new RGB-D capture which we use to update $I^{t}_{base}$ $C^{t}_{base}$ $D^{t}_{base}$ to get $I^{t+1}_{base}$ $C^{t+1}_{base}$ $D^{t+1}_{base}$. Once we have reached our desired stopping criteria, we create and return our final 3D reconstruction $R_{final}$ using $I^{m}_{base}$ $C^{m}_{base}$ $D^{m}_{base}$.
@@
Although we choose to use a simple greedy acquisition strategy, our policy provides the flexibility to use incorporate any desired constraints into the robots’ acquisition strategy. In Section 4, we evaluate VIN-NBV with constraints on the number of acquisitions and time in motion for the robot. Our method also allows the user to specify any specific sampling strategy to create a set of query views to evaluate, providing high adaptability to downstream applications.

### v3

@@ Abstract
Next Best View (NBV) algorithms aim to maximize 3D scene acquisition quality using minimal resources, e.g. number of acquisitions, time taken, or distance traversed. Prior methods often rely on coverage maximization as a proxy for reconstruction quality, but for complex scenes with occlusions and finer details, this is not always sufficient and leads to poor reconstructions. Our key insight is to train an acquisition policy that directly optimizes for reconstruction quality rather than just coverage. To achieve this, we introduce the View Introspection Network (VIN): a lightweight neural network that predicts the Relative Reconstruction Improvement (RRI) of a potential next viewpoint without making any new acquisitions. We use this network to power a simple, yet effective, sequential sampling-based greedy NBV policy. Our approach, VIN-NBV, generalizes to unseen object categories, operates without prior scene knowledge, is adaptable to resource constraints, and can handle occlusions. We show that our RRI fitness criterion leads to a $\sim$30% gain in reconstruction quality over a coverage-based criterion using the same greedy strategy. Furthermore, VIN-NBV also outperforms deep reinforcement learning methods, Scan-RL and GenNBV, by $\sim$40%. For more result visuals please refer to our Project page.
Figure 1: Our approach, VIN-NBV (blue), is an NBV policy that selects best next views by maximizing a predicted reconstruction quality criterion and outperforms coverage maximization policy, Cov-NBV (red), achieving higher reconstruction accuracy with less captures.
@@ 1 Introduction
Acquiring 3D knowledge of an environment is often a crucial step for many robotics applications; e.g., a drone surveying a disaster zone to assist search and rescue efforts, or autonomous robots monitoring large construction and agricultural sites. Real-world environments contain diverse objects of varying sizes, occlusions, and geometries; such complexities currently require a slow dense scan [11, 36, 9, 2, 3, 10, 12, 38] to reconstruct the 3D scene effectively. However, in many search and rescue operations, where speed leads to successful outcomes, acquiring the 3D scene in the shortest time possible is critical. In large construction or agricultural sites, battery life can be a limiting factor, making long captures costly. These challenges motivated techniques for resource-efficient scanning and 3D reconstruction, where constraints such as the number of captures, travel distance, or battery life must be respected. This problem has often been called the Next Best View (NBV) selection problem, where the goal is to predict a set of optimal camera viewpoints to maximize reconstruction quality.
@@
Predicting NBVs is challenging due to the large search space for selecting the pose of each view, efficiently computing the optimal solution in this non-convex space, and the complexity of scene geometry and occlusions. Earlier works on NBV often assume prior knowledge of the scene (e.g., a preliminary scan or CAD model) [7, 39, 45, 42, 21, 46], which limits applications in unexplored environments. Approaches that do not require prior knowledge of the scene predict NBVs by either maximizing coverage in the scene [26, 35, 16, 29, 13, 14] or by maximizing information gain [17, 45, 20] in new viewpoints. While earlier methods have relied on heuristics or optimization [26, 35], recent methods often train a deep reinforcement learning (RL) algorithm [16, 29, 17, 45, 33, 20, 43, 5] to maximize coverage.
@@
While maximizing coverage with RL [5, 29] may lead to a generalizable policy that can provide a good approximation of the scene, it ignores the fact that certain regions from the same scene may have more complicated geometry and self-occlusion. For example, in Fig. 1, if the fence of the house occludes a part of the wall, a single viewpoint capture may satisfy the coverage criterion, but will lead to a poor reconstruction with holes in the wall unless additional viewpoints are captured. Thus, our key idea is to develop an NBV acquisition policy that is trained to directly maximize the 3D reconstruction quality instead of only relying on a coverage criterion, which may fail to revisit areas with low coverage gains that will likely improve the reconstruction quality. Our policy design assumes the agent acquires RGB-D images and has no prior information about the scene.
We introduce a View Introspection Network (VIN) that can predict the relative reconstruction improvement (RRI) of any potential next view over the base images, without actually acquiring any image from the viewpoint. To demonstrate the effectiveness of directly optimizing reconstruction quality over coverage criterion, we design a simple, but effective, greedy sequential next best view selection policy VIN-NBV. At each acquisition step, the policy evaluates a set of uniformly sampled query views around the object, using the VIN to predict their effectiveness in improving reconstruction quality. The view that maximizes reconstruction improvement over base views is chosen. Our proposed VIN-NBV policy is generalizable to unseen object categories, and can operate under varying constraints, e.g. number of acquisitions, time traversed, and distance traversed; it can also avoid collisions.
The VIN is a lightweight neural network that takes the already captured base images, their camera parameters, and the query viewpoint camera parameters as input. The VIN performs a 3D-aware featurization of the already captured base views and then predicts a fitness score for each query view. We use the fitness score to indicate what Relative Reconstruction Improvement (RRI) over base views a query view may have. The VIN is trained using an imitation learning approach, where the RRI of all query viewpoints is pre-computed by explicitly reconstructing the scene with each query view and calculating the reduction in error.
We show that a relative reconstruction improvement (RRI) fitness criterion predicted by the VIN leads to ${\sim}30\%$ improvement in reconstruction quality over a coverage fitness criterion, when using the same sampling-based greedy sequential strategy. We further show that the VIN-NBV policy significantly outperforms existing RL-based algorithms that aim to maximize coverage, reducing reconstruction error by 41% compared to Scan-RL [29] and 39% compared to GenNBV [5].
@@ 2 Related work
NBV methods requiring prior knowledge. Many existing approaches assume prior knowledge of the 3D scene, which limits applications in previously unexplored environments. Some approaches [7, 39, 45] directly exploit a pre-existing 3D model or its approximation. Other methods rely on 2D maps, such as estimating the height of buildings as a rough 3D model [21] or obtaining a 2.5D model of the scene [46]. For scenes lacking pre-existing information, often a drone fly-through along a default trajectory is made to obtain an initial coarse reconstruction [35, 17]. In contrast, our method does not require any prior information about the scene; instead, we start with two adjacent views and sequentially choose the best next views until a specific termination criterion is met.
Optimal view selection with expensive scene representations. A related line of work addresses the problem of selecting an optimal subset of images for reconstruction by using complex scene representations. These representations are often built from existing dense capture. Although the selection algorithms themselves do not require dense pre-capture imagery as input, they rely on scene representations, such as dense point clouds, meshes, or pretrained radiance fields, that are themselves costly to obtain. Earlier approaches targeted optimal view selection for Multi-View Stereo reconstruction [18, 11], while more recent work explored view selection for Neural Radiance Fields (NeRF) and its variants [37, 27, 22, 20]. In contrast, our approach does not assume access to a dense capture or an expensive prebuilt representation; we acquire images directly from the optimal viewpoints predicted by the NBV policy.
Criteria for predicting NBV. Previous approaches use proxy metrics like coverage [26, 29, 1, 16, 5] or information gain [23, 19, 31, 20] for predicting NBVs. State-of-the-art NBV techniques, GenNBV [5] and ScanRL [29], use Reinforcement Learning (RL) with coverage-based reward functions. While coverage-based policies may lead to reasonable reconstruction quality, they fail to account for complex structures and occlusions, often resulting in holes in the reconstructed scene. On the other hand, the information gain criterion is image-based and often lacks 3D knowledge of the scene, restricting generalization. Our work goes beyond coverage by directly maximizing the reconstruction quality, which leads to significant improvement under resource constraints, as shown empirically. Rather than relying on RL [29, 5], we design a greedy sequential policy, VIN-NBV, trained with imitation learning to help predict the reconstruction improvement of next views.
@@ 3 Method
In the following section, we formalize the NBV problem and describe our VIN-NBV policy in detail. We will first provide a mathematical overview of the NBV problem in Sec. 3.1. Then, we will introduce our proposed sampling-based greedy NBV policy, VIN-NBV, in Sec. 3.2, followed by the design of the View Introspection Network (VIN) in 3.3 and its training in 3.4.
@@ 3.1 Problem Setup
Consider an agent that has acquired $k$ initial base images of a scene $I_{base}=\{I_{1},...,I_{k}\}$ from viewpoints with camera parameters $C_{base}=\{C_{1},...,C_{k}\}$ and depth maps $D_{base}=\{D_{1},...,D_{k}\}$ either captured with a depth sensor or predicted with a Monocular Depth Estimator or Multi-View Stereo algorithms. The agent then runs a 3D reconstruction pipeline using the initial base views to reconstruct the 3D scene as $\mathcal{R}_{base}$. The goal of NBV is to predict a set of $m$ next best views $C_{nbv}=\{C_{{1}},...,C_{m}\}$ from which the scene should be captured to maximize the reconstruction quality of the scene.
More specifically, from $C_{nbv}$ we capture NBV images $I_{nbv}=\{I_{1},...,I_{m}\}$ and associated depth maps $D_{nbv}=\{D_{1},...,D_{m}\}$ and perform 3D reconstruction to create $\mathcal{R}_{final}$ using $I_{base}\cup I_{nbv}$ and $D_{base}\cup D_{nbv}$. While previous NBV techniques maximize coverage, we maximize reconstruction quality measured using the relative improvement of the Chamfer Distance of $\mathcal{R}_{final}$ over $\mathcal{R}_{base}$:
$C^{*}_{nbv}=\operatorname*{arg\,max}_{C_{nbv}}~\frac{CD(\mathcal{R}_{base},\mathcal{R}_{GT})-CD(\mathcal{R}_{final},\mathcal{R}_{GT})}{CD(\mathcal{R}_{base},\mathcal{R}_{GT})},$ (1)
@@
Algorithm 1 VIN-NBV Policy
[…]
@@
9:   $\widehat{\mathcal{RRI}}(q_{i})=\text{VIN}_{\theta}(R^{t}_{base},C^{t}_{base},C_{q_{i}})$
@@ 3.2 NBV Policy with View Introspection Network
We propose a simple greedy imitation learning based approach, where for each acquisition, the agent samples a set of query camera viewpoints and evaluates their fitness, and chooses the ‘best’ one and repeats the process until some termination criterion has been reached. This acquisition policy has been described in Algorithm 1.
@@
The VIN-NBV policy begins with $I^{k}_{base}$, $C^{k}_{base}$, $D^{k}_{base}$ then iteratively selects next views until reaching a desired termination criteria, defined to fit downstream task constraints, e.g., number of image acquisitions, time traversed, battery life, etc. At each step, we update the reconstruction $R^{t}_{base}$ by back-projecting the RGB-D captures into 3D space. If only RGB images are available, we reconstruct $R^{t}_{base}$ using Multi-View Stereo or Monocular Depth Estimation. We then sample $n$ query views around this reconstruction to evaluate their fitness and choose the ’best’.
@@
Our key idea is the introduction of the View Introspection Network (VIN), which independently evaluates potential query views and predicts their fitness. Instead of evaluating the fitness of each view to maximize coverage, we focus on maximizing the reconstruction quality. More specifically, we define fitness criterion as the Relative Reconstruction Improvement ($\mathcal{RRI}$) over the base views by capturing the query view $q$ as defined in eq. 2. Relative Improvement is formulated to be independent of object types and scales, which otherwise affects the Chamfer Distance computation.
$\mathcal{RRI}(q_{i})=\frac{CD(\mathcal{R}_{base},\mathcal{R}_{GT})-CD(\mathcal{R}_{base\cup q_{i}},\mathcal{R}_{GT})}{CD(\mathcal{R}_{base},\mathcal{R}_{GT})}.$ (2)
@@
We train VIN to predict the RRI fitness criterion $\widehat{RRI}(q)$ for a query view $q$ by taking the existing reconstruction $R_{base}$, camera parameters $C_{base}$, and query view camera parameters $C_{q}$ as input:
$\widehat{\mathcal{RRI}}(q)=\text{VIN}_{\theta}(R_{base},C_{base},C_{q}).$ (3)
@@
The design of VIN’s neural architecture and its training are described in Sec 3.3 and 3.4. After evaluating all $n$ query views, the VIN-NBV policy greedily selects the one with the highest improvement score. We move to the selected view position $C^{t}_{*}$ and acquire new RGB-D capture which we use to update $I^{t}_{base}$ $C^{t}_{base}$ $D^{t}_{base}$ to get $I^{t+1}_{base}$ $C^{t+1}_{base}$ $D^{t+1}_{base}$. Once we have reached our desired stopping criteria, we create and return our final 3D reconstruction $R_{final}$ using $I^{m}_{base}$ $C^{m}_{base}$ $D^{m}_{base}$.
@@
Although we adopt a simple greedy acquisition strategy, our policy provides the flexibility to incorporate any desired constraints into the agent’s acquisition strategy. In Section 4, we evaluate the VIN-NBV policy with constraints on the number of acquisitions and time in motion for the capture agent. Our method also allows the user to specify any sampling strategy to create a set of query views to evaluate, providing high adaptability to downstream applications.